18.01 PROBLEM SET 8

Hand in to Room 2-108 by 1:00 pm Friday, May 10.

The due date and time are absolute. Late homeworks are accepted until the solutions are posted, but the score will be multiplied by 1/2.

Everything included here may be used later. You do not need to memorize the proofs, but you need to know the facts.

1. SP.VI/1. 337/25. 342/38. 348/3, 17. (5 points)

2. Compute

I = ∫_{0}^{∞} ∫_{x}^{∞} \frac{1 + 2t}{t^{2}(t + 1)^{2}} dtdx.

(15 points)

Typeset by AMS-TEX